### v3

# {cutout}00.3cm22cm1 empty    PointMamba: A Simple State Space Model for Point Cloud Analysis

Abstract
Transformers have become one of the foundational architectures in point cloud analysis tasks due to their excellent global modeling ability. However, the attention mechanism has quadratic complexity and is difficult to extend to long sequence modeling due to limited computational resources and so on. Recently, state space models (SSM), a new family of deep sequence models, have presented great potential for sequence modeling in NLP tasks. In this paper, taking inspiration from the success of SSM in NLP, we propose PointMamba, a framework with global modeling and linear complexity. Specifically, by taking embedded point patches as input, we proposed a reordering strategy to enhance SSM’s global modeling ability by providing a more logical geometric scanning order. The reordered point tokens are then sent to a series of Mamba blocks to causally capture the point cloud structure. Experimental results show our proposed PointMamba outperforms the transformer-based counterparts on different point cloud analysis datasets, while significantly saving about 44.3% parameters and 25% FLOPs, demonstrating the potential option for constructing foundational 3D vision models. We hope our PointMamba can provide a new perspective for point cloud analysis. The code is available at https://github.com/LMD0311/PointMamba.
00footnotetext: Dingkang Liang, Xin Zhou, and Xinyu Wang make equal contribution.
1 Introduction
Point cloud analysis is one of the fundamental tasks in computer vision and has a wide range of real-world applications [48; 31; 66; 5]. It is a challenging task due to the intrinsic irregularity and sparsity of point clouds. To deal with the issues, there has been rapid progress in deep learning-based methods [37; 29; 41; 43], consistently pushing the performance to the new record.
Recently, transformers have shown great potential in point cloud analysis. The key to the transformer is the attention mechanism, which can effectively capture the relationship of a set of points. By further combining self-supervised learning paradigms and fine-tuning the downstream tasks, these transformer-based methods [64; 29; 42] have obtained superior performance. For example, Point-BERT [64] generates discrete point tokens containing meaningful local information from masked tokens. PointMAE [29] follows the pipeline of ViT-based MAE [21], reconstructing the masked point patches. ReCon [42] learn through ensemble distillation from both generative modeling teachers and single/cross-modal contrastive teachers.
Nonetheless, applying full attention mechanisms to long point tokens leads to a significant increase in computational cost demands, a consequence of the attention calculations’ quadratic complexity in both computation and memory. Efforts to enhance attention efficiency often involve limiting the dimensions or stride of the processing windows [62], which, however, inadvertently narrows the extent of receptive fields. This naturally raises a question: how to design a method that operates with linear complexity, thereby retaining the benefits of global receptive fields?
We note the recent advance of the state space models (SSMs) [12; 54; 24]. Specifically, the Structured State Space Sequence Model [15] (S4) has emerged as a promising class of architectures for sequence modeling. Based on S4, Mamba [12] adopt time-varying parameters to the SSM, proposing an efficient hardware-aware algorithm, having global receptive fields with linear complexity. Currently, those SSM-based models mainly focus on long-range and casual data, e.g., language understanding [35], content-based reasoning [12]. Recently, a few concurrent methods [70; 30] pay attention to 2D vision tasks (e.g., classification, segmentation).
Inspired by the successes of SSM in NLP, this paper aims to explore the potential of SSM in point cloud analysis tasks. Through pilot experiments, we find that simply using Mamba [12] can not achieve ideal performance. We argue that the challenge might be the unidirectional modeling of default SSM. In contrast, the self-attention of the transformer is invariant to the permutation and cardinality of the input elements. Thus, in this paper, we introduce the Point State Space Model (denoted as PointMamba), which has global modeling with linear complexity. Specifically, following the standard transformer-based methods [29; 64], we first adopt a simple point tokenizer to generate a series of point tokens. Then, a simple reordering strategy is proposed to scan data along a specific order, enabling the model to causally capture the point cloud structure. Finally, the reordered point tokens will be fed into the mamba encoder to implement global modeling.
Experiments conducted on various point cloud analysis tasks (e.g., classification, part segmentation) demonstrate the effectiveness of our method. Specifically, our PointMamba surpasses the Transformer-based counterpart [29] while having fewer parameters and FLOPs, demonstrating the potential of SSM in the 3D vision tasks.
In summary, our contributions are listed as follows: 1) We propose a simple state space model for point cloud analysis, named PointMamba, which has global receptive fields with linear complexity. We prove that PointMamba is a potential option for constructing the 3D vision foundation models. 2) We introduce a reordering strategy to scan data along a specific order, enabling the PointMamba to causally capture the point cloud structure.
2 Related work
2.1 Point Cloud Transformers
Transformer [51] can capture global input dependencies using the Attention mechanism and has been dominant in NLP [8; 44; 45; 1] and 2D Vision [10; 32; 2; 28]. In 3D vision, the Transformer architecture continues to evolve. Point-BERT [64] and Point-MAE [37] introduce a standard Transformer architecture for self-supervised learning and is applicable to different downstream tasks, as demonstrated in several works [9; 42; 4; 65]. Other researchers [18; 69; 55] mainly concentrate on enhancing Transformers for point clouds. The PCT [18] conducts global attention directly on the point cloud. Point Transformer [69] applies vector attention theory [68] to perform local attention between each point and its adjacent points. The later Point Transformer series [59; 58] further extends the performance and efficiency of the Transformer for different tasks.
Standard transformers can be smoothly integrated into autoencoders using an encoder-decoder design [21], which makes this structure ideal for pre-training and leads to significant performance improvements in downstream tasks [64; 37]. However, the Attention mechanism, a crucial module of the Transformer, has a time complexity of $O(n^{2}d)$, where $n$ represents the length of the input token sequence and $d$ represents the dimension of the Transformer. This implies that as the input sequence grows, the operational efficiency of the Transformer is significantly constrained.
In this work, we propose a State Space Model (SSM) for point cloud analysis without relying on attention while maintaining the sequential, modality-agnostic modeling advantages of the Transformer.
2.2 State Space Models
Drawing inspiration from control theory, linear state space equations have been combined with deep learning to address sequential data modeling. Initiated by a HiPPO matrix [13], LSSL [16] utilizes linear state space equations for sequence data modeling. Subsequently, the Structured State Space Sequence Model [15] (S4) employs a linear state space for contextualization and shows strong performance on various sequence modeling tasks, especially with lengthy sequences. To avoid high computational burden, HTTYH [17], DSS [19], and S4D [14] find that a diagonal matrix in S4 can replace the Diagonal Plus Low Rank structure without compromising performance. SGConv [27] also provides an alternative approach for treating S4 as a global conventional model. To make S4 faster, GSS [35] leverages the gating structure in the gated attention unit to reduce the dimension of the state space module. Most importantly, the S5 layer [49] introduces the parallel scan and the MIMO SSM, enabling the state space model to be efficiently utilized and widely implemented. As ViS4mer [24], S4ND [36], TranS4mer [25] and The Selective S4 model [54] have applied the state space model in the field of computer vision, achieving preliminary success.
Recently, Mamba [12] has achieved a groundbreaking milestone with its linear-time inference and efficient training process by integrating selection mechanism and hardware-aware algorithms into previous works [14; 19; 35]. Building on the success of Mamba, MoE-Mamba [39] combines Mixture of Experts with Mamba to unlock the potential of SSMs for scaling, achieving Transformer-like performance. Operating on bytes, MambaByte [53] outperforms other byte-level models over several datasets and shows competitive results with subword Transformers. Graph-Mamba [52] enhances long-range context modeling in graph networks by integrating a Mamba block with the input-dependent node selection mechanism. For visual application, Vision Mamba [30] and VMamba [70] use bidirectional Vim Block and the Cross-Scan Module respectively to gain data-dependent global visual context. At the same time, U-Mamba [33] and other works [61; 47] show superior performance in medical image segmentation. These are concurrent works that have appeared while this manuscript is being prepared. However, the application of the Mamba for point cloud analysis remains unexplored. In this work, we delve into the potential of Mamba in 3D vision and propose PointMamba.
3 Method
3.1 Preliminaries
State Space Models. State Space Models (SSM) [15] offer a robust framework for modeling physical systems, particularly linear time-invariant (LTI) systems. These models excel in representing systems using a set of first-order differential equations, capturing the dynamics of the system’s state variables:
$\dot{h}(t)$ $=Ah(t)+Bx(t),$ (1)
$y(t)$ $=Ch(t)+Dx(t).$
where $\dot{h}(t)$ denotes the time derivative of the state vector $h(t)$, with matrices $A$, $B$, $C$ and $D$ defining the relationships between the $h(t)$, $x(t)$, $y(t)$.
Since SSMs operate on continuous sequences $x(t)$, they are unable to process discrete token inputs such as images and texts $\{x_{0},x_{1},\ldots\}$. Consequently, the adaptation to a discretized version of the SSM becomes essential:
$h_{k}$ $=\bar{A}h_{k-1}+\bar{B}x_{k},$ (2)
$y_{k}$ $=\bar{C}h_{k}+\bar{D}x_{k}.$
This discretized form of the SSM effectively maps from sequence to sequence and establishes a recursive relationship with respect to the hidden state $h_{k}$. In this framework, $\bar{A}\in\mathbb{R}^{N\times N},\bar{B}\in\mathbb{R}^{N\times 1},\bar{C}\in\mathbb{R}^{N\times 1},\bar{D}\in\mathbb{R}^{N\times 1}$ are all composed of parameter matrices. Notably, $\bar{D}$, typically serving as a residual connection, is often simplified or disregarded in certain contexts. The transition from continuous to discrete form involves discretizing $x(t)$ using a step size $\Delta$, thus framing the input $x_{k}$ as a sample from the underlying continuous signal $x(t)$, specifically $x_{k}=x(k\Delta)$. This leads to the following derivations:
$\bar{A}$ $=(I-\Delta/2\cdot A)^{-1}(I+\Delta/2\cdot A),$ (3)
$\bar{B}$ $=(I-\Delta/2\cdot A)^{-1}\Delta B,$
$\bar{C}$ $=C.$
Selective SSMs. However, the inherent Linear Time-Invariant (LTI) characteristic of SSMs, characterized by the consistent use of matrices $\bar{A},\bar{B}$, $\bar{C}$ and $\Delta$ across various inputs, restricts their ability to filter and understand contextual nuances within diverse input sequences. To address this, we utilize selective SSM as the core operator of our PointMamba. Selective SSMs involve rendering $\bar{B}$, $\bar{C}$, and $\Delta$ as dynamic, input-oriented elements, effectively transitioning the SSM into a time-variant model. This change allows the model to adapt more effectively to different input contexts, enhancing its capacity to capture relevant temporal features and relationships, thereby leading to a more accurate and efficient representation of the input sequence.
Figure 1: The pipeline of our PointMamba. We start by using FPS and KNN to obtain point patches, and then embed them into point tokens. After that, we reorder and triple the sequence. Finally, we feed the token sequence into the Mamba blocks.
3.2 PointMamba
3.2.1 Overall
The pipeline of our method is shown in Fig. 1, consisting of the point tokenizer, a reordering strategy, a series of mamba blocks, and a downstream head. Starting with an input point cloud, we employ a lightweight PointNet to embed point patches and generate point tokens. Subsequently, a simple yet effective reordering strategy is applied to organize point tokens according to geometric coordinates and triple the sequence. Finally, the entire sequence is fed to Mamba blocks.
3.2.2 Point Tokenizer
Building on the pioneer works [64; 37], we utilize the Farthest Point Sampling (FPS) and K-Nearest Neighborhood (KNN) algorithm to partition the input point cloud into irregular point patches. Specifically, given an input point cloud with $M$ points $I\in\mathbb{R}^{M\times 3}$, the FPS is applied to sample $n$ key points from the original point cloud $I$. Subsequently, the KNN algorithm selects $k$ nearest neighbors for each key point, forming $n$ patches $P\in\mathbb{R}^{n\times k\times 3}$ with patch size $k$. To aggregate local information, points within each patch are normalized by subtracting the key point to obtain relative coordinates. We finally map the unbiased local patches to feature space using a lightweight PointNet [40], obtaining point tokens $E_{0}\in\mathbb{R}^{n\times C}$.
3.2.3 Reordering Strategy
Despite its distinctive features, Mamba processes input data causally (i.e., unidirectional modeling), limiting its ability to capture information beyond the scanned portion. This makes it well-suited for 1-D sequence modeling, such as NLP tasks but presents challenges when working with non-causal data. For example, the patches of a point cloud sample are random and unordered.
Figure 2: A simple illustration of the reordering strategy. The point patches are reordered and concatenated along the $x$, $y$, and $z$ axes.
A simple approach is to scan data along a specific order, enabling the model to causally capture the structure of the point cloud. As shown in Fig. 2, the original random sequence used in transformer-based counterparts [64; 37] lacks a coherent geometric order, making it challenging for unidirectional modeling Mamba. To address this, we arrange point tokens $E_{0}\in\mathbb{R}^{n\times C}$ based on the geometric coordinates of its clustering center, sorting and concatenating along the $x$, $y$, and $z$ axes, thereby tripling the length of the point token (i.e., $E_{0}^{\prime}\in\mathbb{R}^{3n\times C}$). We argue that applying the reordering strategy improves Mamba’s global modeling capability by providing a more logical geometric scanning order. We find that despite increasing the sequence length, the reordering strategy results in a considerable performance improvement without a significant increase in FLOPs.
3.2.4 Mamba Block
After obtaining the reordered token $E_{0}^{\prime}$, we will feed it into the encoder, containing a series of mamba blocks. Specifically, for each mamba block, layer normalization (LN), SSM, depth-wise convolution [7], and residual connections are employed. A stand mamba layer is shown in Fig. 1, and the output can be summarized as follows:
$Z^{\prime}_{l}$ $=DW(MLP(LN(Z_{l-1}))),$ (4)
$Z_{l}$ $=MLP(LN(SSM(\sigma(Z^{\prime}_{l})))\times\sigma(LN(Z_{l-1})))+Z_{l-1},$
where $Z_{l}\in\mathbb{R}^{3n\times C}$ is the output of the $l$-th block and $Z_{0}$ is equal to $E^{{}^{\prime}}_{0}$. DW means the depth-wise convolution. Following the DW, a SiLU [22] and SSM are adopted. The SSM is the key to the mamba block, with a detailed description listed in Sec. 3.1.
3.2.5 Pre-training
For pre-training, we follow the Point-MAE [37] settings. Firstly, we randomly mask the point patches with a high ratio of 60%. Then, an asymmetric autoencoder is employed to extract the point feature, and the final layer of the autoencoder utilizes a simple prediction head for reconstruction.
Specifically, the autoencoder can be summarized as Eq. 5. The encoder $\mathcal{F}_{e}$ takes the visible token $T_{v}$ (unmasked) as input, while the decoder $\mathcal{F}_{d}$ takes the encoder’s outputs and the learnable masked token $T_{m}$ as input. Unlike Point-MAE, we only add the positional embedding ($PE$) to the first layer of both the encoder and the decoder. One key aspect of our Mamba decoder is that it models sequential data, enabling it to infer the masked information from the unmasked token.
$T_{v}$ $=\mathcal{F}_{e}(T_{v}+PE),$ (5)
$H_{v},H_{m}$ $=\mathcal{F}_{d}\left(concat\left(T_{v},T_{m}\right)\right),$
$P_{m}$ $=\mathcal{F}_{h}(H_{m})$
To reconstruct masked point patches in coordinate space, we employ a Linear head $\mathcal{F}_{h}$ to project the masked token $H_{m}$ to the vector $P_{m}$ with the same shape as the masked input points. The Chamfer Distance [11] is then used as the reconstruction loss to recover the coordinates of the points in each masked point patch.
4 Experiments
In this section, we first present the implementation details. Then, we evaluate the proposed model with three downstream tasks, including synthetic object classification, real-world object classification and part segmentation. We also carry on an ablation study for our PointMamba.
4.1 Implementation Details
Table 1: Implementation details for pre-training and downstream tasks.
Configuration Pre-training Classification Segmentation
ShapeNetCore ModelNet40 ScanObjectNN ShapeNetPart
Optimizer AdamW AdamW AdamW
Learning rate 1e-3 3e-4 5e-4 2e-4
Weight decay 5e-2 5e-2 5e-2 5e-2
Learning rate scheduler cosine cosine cosine cosine
Training epochs 300 300 300 300
Warmup epochs 10 10 10 10
Batch size 128 32 32 16
Num. of encoder layers $N$ 12 12 12 12
Num. of decoder layers 4 - - -
Input points $M$ 1024 1024 2048 2048
Num. of patches $n$ 64 64 128 128
Patch size $k$ 32 32 32 32
Augmentation Scale&Trans Scale&Trans Scale&Trans -
SSM is a newcomer to 3D point cloud analysis, and no existing works currently detail the specific implementation. For pre-training, we follow Point-MAE [37], using the ShapeNetCore training set, a subset of ShapeNet [3]. When presenting the results for downstream tasks, we include both scratch and pre-training scenarios.
For classification tasks, we do not use the classification token [CLS] as PointMamba can only perform unidirectional modeling, unlike ViT[10], which utilizes an attention mechanism for global interactions. Instead, we simply select the average pooling of the last block’s output for classification. For part segmentation, our segmentation head follows Point-MAE [37] and leverages features from the 3-rd, 7-th, and 11-th layers of the Mamba block for fair comparisons. We merge the three feature levels, apply max and average pooling for global features, and then concatenate per-point features with the global ones. Finally, a Linear layer is used to forecast the label for each point.
All experiments are carried out using a single RTX 4090 GPU. More implementation details can be found in Tab. 1.
Table 2: Object classification on ModelNet40 [60]. Overall accuracy (%) is reported. denotes reproduced results. The results obtained from inputting 1024 points without voting are reported.
Methods Backbone Param. (M) $\downarrow$ FLOPs (G) $\downarrow$ Points Num. OA (%) $\uparrow$
PointNet [40] - 3.5 0.5 1k 89.2
PointNet++ [41] - 1.5 1.7 1k 90.7
PointCNN [26] - 0.6 - 1k 92.2
DGCNN [38] - 1.8 2.4 1k 92.9
PointNeXt [43] - 1.4 1.6 1k 92.9
PCT [18] Transformer-based 2.9 2.3 1k 93.2
OctFormer [55] Transformer-based - - 1k 92.7
Training from scratch
Point-MAE [37] Transformer-based 22.1 2.4 1k 92.3
PointMamba (ours) Mamba-based 12.3 1.8 1k 92.4
Training from pre-training
Point-BERT [64] Transformer-based 22.1 2.4 1k 92.7
Point-MAE [37] Transformer-based 22.1 2.4 1k 93.2
PointMamba (ours) Mamba-based 12.3 1.8 1k 93.6
Table 3: Object classification on ScanObjectNN [50]. We evaluate PointMamba on three variants, with PB-T50-RS being the most challenging. Accuracy (%) is reported. denotes reproduced results, indicates that using simple rotational augmentation [9] for training.
Methods Backbone Param. (M) $\downarrow$ FLOPs (G) $\downarrow$ OBJ-BG $\uparrow$ OBJ-ONLY $\uparrow$ PB-T50-RS $\uparrow$
PointNet [40] - 3.5 0.5 73.3 79.2 68.0
PointNet++ [41] - 1.5 1.7 82.3 84.3 77.9
PointCNN [26] - 0.6 0.9 86.1 85.5 78.5
DGCNN [56] - 1.8 2.4 82.8 86.2 78.1
PRANet [6] - - - - - 81.0
MVTN [20] - 11.2 43.7 - - 82.8
PointNeXt [43] - 1.4 1.6 - - 87.7
PointMLP [34] - 13.2 31.4 - - 85.4
RepSurf-U [46] - 1.5 0.8 - - 84.3
ADS [23] - - - - - 87.5
Training from scratch
Transformer [64] Transformer-based 22.1 4.8 79.86 80.55 77.24
Point-MAE [37] Transformer-based 22.1 4.8 86.75 86.92 80.78
PointMamba (ours) Mamba-based 12.3 3.6 88.30 87.78 82.48
Training from pre-training
Point-BERT [64] Transformer-based 22.1 4.8 87.43 88.12 83.07
Point-MAE [37] Transformer-based 22.1 4.8 90.02 88.29 85.18
Point-MAE [37] Transformer-based 22.1 4.8 92.77 91.22 89.04
PointMamba (ours) Mamba-based 12.3 3.6 90.71 88.47 84.87
PointMamba (ours) Mamba-based 12.3 3.6 93.29 91.91 88.17
Table 4: Part segmentation on the ShapeNetPart [63]. The mIoU for all classes (Cls.) and for all instances (Inst.) are reported. denotes reproduced results.
Methods Backbone Param. (M) $\downarrow$ FLOPs (G) $\downarrow$ Cls. mIoU (%) $\uparrow$ Inst. mIoU (%) $\uparrow$
PointNet [40] - - - 80.39 83.7
PointNet++ [41] - - - 81.85 85.1
DGCNN [56] - - - 82.33 85.2
APES [57] - - - 83.67 85.8
Training from scratch
Point-MAE [37] Transformer-based 27.1 15.5 83.91 85.7
PointMamba (ours) Mamba-based 17.4 14.3 83.94 85.8
Training from pre-training
Point-MAE [37] Transformer-based 27.1 15.5 84.19 86.1
PointMamba (ours) Mamba-based 17.4 14.3 84.42 86.0
4.2 Compared with Transformer-based Counterpart
Synthetic Object Classification on ModelNet40. ModelNet40 [60] is a pristine 3D CAD dataset comprising 40 categories. As shown in Tab. 2, we report overall Accuracy without voting. PointMamba achieves competitive results compared to Point-MAE from scratch while using fewer parameters and FLOPs. Using the MAE-like pre-training strategy, our PointMamba achieved a 1.2% performance improvement, outperforming Point-BERT and Point-MAE by 0.9% and 0.4%, respectively. The pre-trained PointMamba can also achieve competitive results against some point cloud-specific transformer models such as PCT and OctFormer.
Real-World Object Classification on ScanObjectNN. ScanObjectNN [50] is a highly challenging 3D dataset consisting of about 15,000 objects across 15 categories. These objects are scanned from real-world indoor scenes with cluttered backgrounds, adding to the complexity of object classification. In Tab. 3, we conduct experiments on three versions of ScanObjectNN (i.e., OBJ-BG, OBJ-ONLY, and PB-T50-RS) trained from scratch and pre-training, each with increasing complexity. It is worth noting that our PointMamba outperforms the Point-MAE in all three variants when training from scratch while reducing 44.3% parameters and 25% FLOPs. Specifically, compared with reproduced results, PointMamba achieves increases of 1.55%, 0.86%, and 1.7% in the three ScanObjectNN variants. Even when trained from pre-training, our model still achieves competitive results. Specifically, the PointMamba outperforms Point-BERT 3.28%, 0.35% and 1.18% on three variants, exceeding Point-MAE 0.69% and 0.18% on OBJ-BG and OBJ-ONLY. Further adopting simple rotational augmentation [9] can further improve the model performance, showcasing the great potential for real-world classification. Another thing to notice is that, with the same pre-training strategy, the improvement of our PointMamba is not as significant as Point-MAE but still surpasses Point-MAE in most cases, suggesting that a customized pre-training strategy for PointMamba is a worthwhile research direction, which left our future work.
Figure 3: Qualitative results of PointMamba and Point-MAE on ShapeNetPart.
Part Segmentation on ShapeNetPart. Part segmentation on ShapeNetPart [63] is another challenging task aiming to predict a more detailed label for each point within a sample. We collect 2048 points in 128 point patches, then adopt the reordering strategy to triple the input. As shown in Tab. 4, we report mean IoU (mIoU) for all classes (Cls.) and all instances (Inst.). Our PointMamba achieve competitive results while significantly reducing parameters. Note that the increase in parameters and FLOPs in segmentation mainly comes from the head part. We do not use the U-Net structure for fair comparison, which may limit the performance. Moreover, we present qualitative results of our PointMamba and Point-MAE in Fig. 3. It can be found that PointMamba achieves highly competitive results to the Transformer-base counterpart [29].
Figure 4: Comparisons of inference GPU memory.
GPU Memory with Lengthy Sequences. Previous research [67] reveals that vanilla vision transformer (ViT) faces challenges in processing large-scale point clouds due to GPU memory constraints. As shown in Fig. 4, we evaluate the inference memory usage of  PointMamba and Point-MAE on lengthy sequences with batch size 32 on a single NVIDIA A800. The results illustrate that the memory efficiency of PointMamba offers a promising prospect for addressing large-scale point clouds.
4.3 Ablation study
To investigate the architecture design, we conduct ablation studies on ScanObjectNN [50] with training from scratch. The default setting for PointMamba is marked in green.
Table 5: The effect of reordering strategy. denotes reproduced results.
Methods Param. (M) FLOPs (G) OBJ-BG OBJ-ONLY
Point-MAE (1x) 22.1 4.8 86.75 86.92
Point-MAE (3x) 22.1 10.2 86.40 86.57
Point-MAE (6x) 22.1 18.4 86.06 86.23
PointMamba (1x) 12.3 2.6 87.61 86.40
PointMamba (3x) 12.3 3.6 88.30 87.78
PointMamba (6x) 12.3 5.2 87.44 88.30
Analysis on reordering strategy. The reordering strategy adapts PointMamba to random and unordered point clouds data. Given input point tokens $E_{0}\in\mathbb{R}^{n\times C}$, the sequence is copied, sorted, and tripled in length to yield $E_{0}^{\prime}\in\mathbb{R}^{3n\times C}$. We conduct experiments on training from scratch to prove the effectiveness of the proposed reordering strategy in Tab. 5, where ‘1x’ denotes no use of the reordering strategy, ‘3x’ denotes our default reordering operation, and ‘6x’ denotes the concatenation of the flipped ‘3x’ sequence, resulting in a sequence six times longer. It is observed that: (i) The Transformer’s global interaction through the attention mechanism results in a slight drop when the sequence length grows. (ii) Reordering strategy does help unidirectional modeling PointMamba adapt to non-causal point cloud data, bring 0.69% and 1.38% increase in two variants. (iii) Further increasing the sequence length can improve the performs, we apply ‘3x’ by considering the trade-off between accuracy and FLOPs. (iv) While tripling the sequence, our PointMamba only has a slight increase in FLOPs compared to Point-MAE (1.0G vs. 5.4G), suggesting that PointMamba is highly competitive as its linear complexity. PointMamba utilizes 1D sequence modeling to incorporate global information in the final hidden state.
Table 6: The effect of classification token $[\mathtt{CLS}]$.
Methods Param. (M) FLOPs (G) OBJ-BG OBJ-ONLY
Before the sequence 12.4 3.6 88.12 87.09
After the sequence 12.4 3.6 87.44 86.75
No $[\mathtt{CLS}]$ 12.3 3.6 88.30 87.78
Analysis on classification token. Previous works [10; 64; 37] often use a classification token $[\mathtt{CLS}]$ as a global token for classification. As in Tab. 6, we find that without $[\mathtt{CLS}]$ and utilizing only the average pooling of the final block’s output yields the best results for PointMamba.
Table 7: The effect of dimension $C$.
Dimension $C$ Param. (M) FLOPs (G) OBJ-BG OBJ-ONLY PB-T50-RS
128 2.1 2.2 87.78 87.44 81.61
256 6.0 2.8 88.81 87.61 81.16
384 12.3 3.6 88.30 87.78 82.48
512 21.3 4.8 88.47 87.61 82.13
Analysis on dimension $C$. The dimension $C$ for the Mamba block is a crucial hyper-parameter for controlling the model’s scale. As in Tab. 7, we empirically set $C$ to 384 by considering the trade-off in performance across three variants.
4.4 Limitation
The main limitations of PointMamba include the need to triple the sequence in the reordering strategy, which lacks elegance and slightly reduces the capability to model very long sequences. Besides, the pre-training strategy is just a preliminary attempt without considering the unidirectional modeling characteristic of Mamba, which is also part of our future work.
5 Conclusion
In this study, we propose a state space model (SSM) for point cloud analysis named PointMamba. One challenge with applying unidirectional SSM is that it cannot capture non-causal data like point clouds. By taking embedded point patches as input and applying a simple reordering strategy, we adapt point cloud data for the unidirectional data-dependent SSM, i.e., Mamba. We also attempted to pre-train PointMamba, further expanding the potential of our method. Experimental results show that PointMamba outperforms the transformer-based counterparts on different point cloud analysis datasets while significantly reducing parameters and FLOPs. As a newcomer to point cloud analysis, PointMamba is a promising option for constructing 3D vision foundation models, and we hope it can offer a new perspective for the field.